OpenStudy (anonymous):

how would i find the vert. and horz. asymtope of y=5 over x?

OpenStudy (anonymous):

vertical asymptote is where the equation is undefined. \[y=\frac{1}{x}\] is not defined at 0 because \[\frac{1}{0}\] is not a number. so the vertical asymptote is given by the equation \[x=0\] otherwise known as the y- axis. the horizontal asymptote is the line the graph approaches as \[x->\infty\] as x gets large \[\frac{1}{x}\] gets close to zero. think about \[\frac{1}{1000000}=.000001\] likewise \[x->-\infty\] y also gets very small. so the horizontal asymptote is the line \[y=0\]

OpenStudy (anonymous):

find vertical asymptote: set denominator = 0 and solve. horizontal asymptote: a) if the degree of the denominator is greater than the degree of the numerator it is y - 0 b) if the degree of the numerator is bigger there isn't one. c) if the degrees are the same it is y = ratio of leading coefficients.